OpenStudy (anonymous):

If the cat takes 43 minutes to complete the 130m displacement and 15 minutes to complete the 84m displacement, what are the magnitude and direction of its average velocity during this 58-minute period of time?

OpenStudy (shamim):

tell me 43+15=?

OpenStudy (anonymous):

Thats 58.

OpenStudy (anonymous):

Would I take the 130/43 & 84/15 to get magnitude?

OpenStudy (shamim):

130+84=214m

OpenStudy (shamim):

so total time is 58 min nd total displacement is 214m

OpenStudy (shamim):

we know average velocity v=total displacement/ total time

OpenStudy (shamim):

average velocity=214/58=?

OpenStudy (anonymous):

average velocity=3.79

OpenStudy (anonymous):

This may help if you need it. I have one try left on the first problem so gotta get it right the first time. Thanks.

OpenStudy (anonymous):

I don't know what to do with the average velocity. Do I take the 130/43 & 84/15 and divide both by 2 to get the magnitude?